# Flexible Music-Conditioned Dance Generation with Style Description Prompts

### v1

@@ IV-C Flexible Generation with Masking
where the mask $M\in{\{0,1\}}^{F\times J}$ can be changed arbitrarily with 0 or 1 values in both the temporal and spatial dimensions, $F$ and $J$ represent the number of frames and joints, respectively. This process is iterated by taking ${\hat{\sigma}}_{t-1}$ as the noise for the next iteration. The spatial-temporal mask supports any combination of temporal constraints and joint constraints. This masking framework offers a robust tool for subsequent applications to generate dance sequences that precisely adhere to arbitrary constraints. Details of the algorithm are shown in Algorithm 2.
Require : The diffusion timestep $T$, music conditioning $c$, style description prompts $s$ and the trained dance generation network $\hat{x}_{\theta}$.
Input : The padded dance sequences $x_{0}$, corresponding masks $M\in{\{0,1\}}^{F\times J}$.
Output : The completed dance sequences $x$.
Randomly sample a noise $\hat{\sigma}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$;
for $t=T,T-1,...1$ do
       Randomly sample a noise $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$ if $t>1$, else $\epsilon=0$.
      For the known sequences $x_{0}$, add noise on them with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t-1}$: $x^{known}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}x_{0}+(1-\overline{\alpha}_{t-1})\epsilon$.
      For the unknown sequences, first use pre-trained network $\hat{x}_{\theta}$ to reconstruct $\hat{x_{0}}$ and then add noise on it with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t-1}$: $x^{unknown}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}\hat{x}_{\theta}(\hat{\sigma}_{t},t,c,s)+(1-\overline{\alpha}_{t-1})\epsilon$.
      Combine $x^{known}_{t-1}$ and $x^{unknown}_{t-1}$ with a spatial-temporal mask $M$: ${\hat{\sigma}}_{t-1}=M\odot x^{known}_{t-1}+(1-M)\odot x^{unknown}_{t-1}$.
end for
Algorithm 2 Flexible Dance Generation Algorithm with Masking
@@
The proposed framework can also be applied for long-term dance generation. Since the model generates a set of frames of a dance sequence at once, increasing the maximum sequence length results in a linear increase in computational cost. In addition, dance generation requires the music condition to match the motion sequence in length, which further extends the memory requirements. We follow the long-term generation strategy [6] to achieve temporal consistency between multiple sequences in long-term generation.
V Experiments
V-A Dataset and Implementation Details
@@
The AIST++ dataset [5] includes 1,408 high-quality dance movements synchronized with music from 10 genres. It comprises motion sequences of varying durations, ranging from 7 seconds to 50 seconds, with an average duration of 13 seconds. We follow the original train/test splits. Consistent with EDGE [6], instances in the training set are cut to 5 seconds at 30 FPS with a stride of 0.5 seconds, and instances in the testing set are cut to 5 seconds at 30 FPS with a stride of 2.5 seconds.
The proposed model has 49.7 million (M) parameters, and was trained on one NVIDIA RTX A6000 GPU for 2 days with a batch size of 128. The implementation is based on the PyTorch. The number of iterations is 1000. The learning rate is 0.0002 and the weight decay is 0.02. The music condition transformer encoder has a depth of 2 layers, with 8 attention heads and a dimension of 512. The decoder block has a depth of 8 layers, with 8 attention heads and a dimension of 512. For long-term generation, we use linear weighted summation to sum up corresponding parts of different slices (5s). Specifically, we perform a weighted summation of the latter half (2.5s) of the previous slice and the former half (2.5s) of the next slice. To mitigate the impact of randomness of stochastic, all evaluation metrics obtained by averaging over 100 trials.
Following [19, 6], we adopt the following evaluation metrics:
@@
Beat Alignment Score: It is used to measure the alignment between the music and the generated movement. We calculate the average time distance between each music beat and its closest dance beat as the beat alignment score. The music beats are extracted using the Librosa library [45], and the kinematic beats are calculated based on the local minima of the velocity of motion joints.
Physical Foot Contact Score: It is a quantitative acceleration-based metric used to score the physical plausibility of the generated kinematic motion. The metric derives from two simple and related observations: (1) In the horizontal plane, any centre of mass (COM) acceleration must be due to static contact between the feet and the ground. Thus, either at least one foot is stationary on the ground or there is no acceleration of COM. (2) On the vertical axis, any positive COM acceleration must be due to static foot contact.
@@
Frechet Inception Distance(FID): FID is the distance between the generated dance movement distribution and the real movement distribution. We calculate the FIDs between the generated dances and the AIST++ dataset for all motion sequences in the kinetic feature space [46] (denoted as $FID_{k}$) and geometric feature space [5] (denoted as $FID_{g}$).
@@
Diversity: This metric assesses the diversity of dance motions by computing the average Euclidean distance between different generated dance motions. Similarly, we compute the Diversity between all motion sequences of the generated dances in the kinetic feature space [46] (denoted as $Div_{k}$) and the geometric feature space [5] (denoted as $Div_{g}$).
@@
TABLE I: Comparision of dance generation on the AIST++ dataset. For simplicity, $\uparrow$ means higher is better, $\downarrow$ means lower is better and $\rightarrow$ means closer to the ground truth is better. During inference, akin to EDGE [6], the guidance weight $w$ is set to 2 by default, and the results obtained with $w=1$ are also reported to explore the influence of amplified conditions.
@@
Method Guidance weight $w$ = 2 Guidance weight $w$ = 1 $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ EDGE [6] 0.26 1.56 35.35 18.92 5.32 4.90 0.25 1.19 45.41 19.42 4.82 5.24 DGSDP (ours) 0.31 1.51 34.32 19.04 6.29 5.81 0.28 1.65 37.21 18.19 7.25 6.42 Ground Truth 0.35 1.33 9.43 7.33 0.35 1.33 9.43 7.33
TABLE II: Results of long-term dance generation on the AIST++ dataset.
@@
Method 7.5 seconds 10 seconds $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ EDGE [6] 0.26 1.11 59.43 25.60 2.97 3.52 0.25 0.88 68.63 32.24 2.55 3.14 DGSDP (ours) 0.30 1.38 55.65 20.74 4.10 5.74 0.30 2.13 56.15 29.98 5.21 6.45 Ground Truth 0.38 1.04 9.34 7.47 0.49 1.70 9.34 7.47
V-B Experimental Settings
We evaluate the performance of flexible dance generation in the following scenarios.
Dance Generation: These setting generates dance movements of 5-second length given the conditioned music. This is the standard setting for music-conditioned dance generation.
Long-Term Dance Generation: This setting aims to generate 7.5-second and 10-second dance movements. For fair comparison, the model is trained on 5-second clips.
@@ V-B Experimental Settings
Seed Motion: Given the real motions of the first and second frames, the model generates dance movements of the following frames.
@@
In-betweening: Given the real motions of the first and last frames, the model completes the motions of the in-between frames.
@@
V-C Results of Dance Generation
We choose EDGE [6] as our baseline. For different approaches, we generate 20 pieces of dances with 5-second length, covering 10 different genres of AIST++. Quantitative results are summarized in Table I. Our method performs favorably against EDGE [6] on most metrics. Specifically, there is a significant improvement in the Beat Alignment Score metric, which is main metric of dance generation, suggesting that the dances generated of our method are more in tune with the musical melody. Moreover, our method achieves the lowest $FID_{k}$, which is defined on motion velocities and energies, reflecting the physical characteristics of dance. The superiority of $FID_{k}$ demonstrates that the proposed DGSDP can generated more realistic dance movements. In addition, our method achieves the closest approximation to the ground-truth values of $Div_{k}$ and $Div_{g}$, revealing the capability to generate more diverse dance movements of our method rather than converging to a few fixed future motions.
V-D Results of Long-Term Dance Generation
Since both the training and test sets are composed of five-second slices, in order to validate the ability of our model to generate long-term dance sequences, we generate the corresponding dances for music that are 7.5 seconds and 10 seconds. We select the pieces of dances that are longer than 7.5 seconds or 10 seconds for long-term dance generation. Experimental result of long-term dance generation on AIST++ is shown in Table II.
Our method performs significantly better than EDGE [6] on most metrics. Specifically, our model achieves a higher Beat Alignment Score, maintaining good performance on the correlation between dance movements and music. What’s more, our model performs better on EDGE [6] on metrics $FID_{k}$,$FID_{g}$,$Div_{k}$ and $Div_{g}$. This reveals that DGSDP can also generate dance movements with better quality and better diversity in long-term dance generation tasks. It is worth noting that the performance of our model on $Div_{k}$ and $Div_{g}$ do not decrease significantly over time like EDGE [6]. This suggests that, despite for long music, the dance movements generated using our method can ensure diversity of movements, avoiding repetitive movements.
V-E Results of Flexible Dance Generation
With the spatial-temporal masking strategy, our approach is applicable to generate dance movements according to various scenarios. We quantitatively compare results of ours with those of EDGE [6] for flexible dance generation in Table III.
In the seed motion scenario and upper-body generation scenario, our method performs favorably against EDGE [6] on all metrics. In the other three scenarios, our method performs better than EDGE [6] on most metrics. This suggests that our approach has a strong capacity for flexible generation from both the temporal and spatial perspectives. Furthermore, it is clearly found that the performance of our model on Beat Alignment Score is close to the ground truth or even exceeds it, revealing the potential for strong self-improvement properties of our model in harmonising music and movement. Overall, our approach is suitable for a wide range of flexible scenarios, and with the help of known information, our model may perform better.
@@ V-E Results of Flexible Dance Generation
TABLE III: Results of flexible dance generation.
@@
Scenarios Method $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$
Seed motion EDGE [6] 0.27 1.19 32.62 21.68 5.81 5.15
DGSDP (ours) 0.33 1.18 27.05 17.36 7.34 5.79
In-betweening EDGE [6] 0.28 1.17 37.33 15.16 4.58 4.86
DGSDP (ours) 0.34 1.24 30.20 14.75 5.80 6.16
Inpainting EDGE [6] 0.36 2.15 21.57 17.50 7.47 8.36
DGSDP (ours) 0.37 1.71 23.80 18.78 7.03 8.34
Upper-body generation EDGE [6] 0.29 1.67 32.90 14.82 5.05 5.45
DGSDP (ours) 0.33 1.40 32.86 14.36 5.39 6.11
Lower-body generation EDGE [6] 0.33 0.92 39.81 15.44 5.28 6.86
DGSDP (ours) 0.36 0.99 42.26 14.73 5.39 7.37
Ground Truth 0.35 1.33 9.43 7.33
[…]

### v2

@@ III-C Spatial-Temporal Masking
where the mask $M\in{\{0,1\}}^{F\times J}$ can be changed arbitrarily with 0 or 1 values in both the temporal and spatial dimensions, $F$ and $J$ represent the number of frames and joints, respectively. This process is iterated by taking ${\hat{\sigma}}_{t-1}$ as the noise for the next iteration. The spatial-temporal mask supports any combination of temporal constraints and joint constraints. This masking strategy offers a robust tool for subsequent applications to generate dance sequences that precisely adhere to arbitrary constraints.
@@
The spatial-temporal masking can also be applied for long-term dance generation. Since the model generates a set of frames of a dance sequence at once, increasing the maximum sequence length results in a linear increase in computational cost. In addition, dance generation requires the music condition to match the motion sequence in length, which further extends the memory requirements. We follow the strategy [9] to achieve temporal consistency between multiple sequences in long-term dance generation.
III-D Prompts of Dance Style
Incorporating dance style is essential for producing high-quality, emotional and contextually appropriate dance movements. For the Style-Guided Motion Diffusion, we explore three encoding methods for prompts of dance style.
Genre Name: To obtain a more semantic representation, we utilize CLIP [35] to extract features for the words associated with the music genre. This embedding captures deep contextual relationships within the textual genre, resulting in more accurate feature representations.
Style Description: Large-scale language models pre-trained on massive text [36] demonstrate impressive capabilities for text generation tasks [37]. To get a more detailed characteristics of the music style, we use GPT-3 [38] to generate style description for dance. We choose the following function to acquire the style description prompts: ”Please generate a detailed description of the dance [$g$], including the characteristics of the dance in terms of body movement.”, where $g$ is the dance genre.
IV Experiments
IV-A Dataset and Evaluation Metrics
@@
The AIST++ dataset [6] includes 1,408 high-quality dance movements synchronized with music from 10 genres. It comprises motion sequences of varying durations, ranging from 7 seconds to 50 seconds, with an average duration of 13 seconds. We follow the train/test splits of EDGE [9]. Instances in the training set are cut to 5 seconds at 30 FPS with a stride of 0.5 seconds, and instances in the testing set are cut to 5 seconds at 30 FPS with a stride of 2.5 seconds.
The Beat Alignment Score ($BeatAlign$), Physical Foot Contact Score ($PFC$), Frechet Inception Distance ($FID$) and Diversity ($Div$) are used as evaluation metrics.
@@
Beat Alignment Score: It is used to measure the alignment between the music and the generated movement. We calculate the average time distance between each music beat and its closest dance beat as the beat alignment score.
Physical Foot Contact Score: It is a quantitative acceleration-based metric used to score the physical plausibility of the generated kinematic motion.
@@
Frechet Inception Distance(FID): FID is the distance between the generated dance movement distribution and the real movement distribution. We calculate the FIDs between the generated dances and the AIST++ dataset for all motion sequences in the kinetic feature space ($FID_{k}$) and geometric feature space ($FID_{g}$).
@@
Diversity: This metric assesses the diversity of dance motions by computing the average Euclidean distance between different generated dance motions. Similarly, we compute the Diversity between all motion sequences of the generated dances in the kinetic feature space ($Div_{k}$) and the geometric feature space ($Div_{g}$).
IV-B Implementation Details
The proposed model has 49.7 million (M) parameters, and was trained on one NVIDIA RTX A6000 GPU for 2 days with a batch size of 128. The implementation is based on the PyTorch. The number of iterations is 1000. The learning rate is 0.0002 and the weight decay is 0.02. The music condition transformer encoder has a depth of 2 layers, with 8 attention heads and a dimension of 512. The decoder block has a depth of 8 layers, with 8 attention heads and a dimension of 512. For long-term generation, we use linear weighted summation to sum up corresponding parts of different slices (5s). We perform a weighted summation of the latter half (2.5s) of the previous slice and the former half (2.5s) of the next slice. To mitigate the impact of randomness of stochastic, all evaluation metrics obtained by averaging over 100 trials.
@@ IV-B Implementation Details
TABLE I: Comparision of dance generation on the AIST++ dataset. For simplicity, $\uparrow$ means higher is better, $\downarrow$ means lower is better and $\rightarrow$ means closer to the ground truth is better. During inference, akin to EDGE [9], the guidance weight $w$ is set to 1 by default, and the results obtained with $w=2$ are also reported to explore the influence of amplified conditions.
@@
Method Model Guidance weight $w$ = 2 Guidance weight $w$ = 1 $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ FACT [6] Others 0.20 30.39 561.14 170.36 Bailando [15] VQ-VAE 0.21 1.72 24.30 20.81 6.83 7.69 TM2D [16] VQ-VAE 0.19 3.28 15.37 28.35 9.10 7.91 EDGE [9] Diffusion 0.26 1.56 35.35 18.92 5.32 4.90 0.25 1.19 45.41 19.42 4.82 5.24 Lodge [11] Diffusion 0.24 37.09 18.79 5.58 4.85 SGMD (ours) Diffusion 0.31 1.51 34.32 19.04 6.29 5.81 0.28 1.65 37.21 18.19 7.25 6.42 Ground Truth 0.35 1.33 9.43 7.33 0.35 1.33 9.43 7.33
TABLE II: Results of long-term dance generation on the AIST++ dataset. We note that very few works focus on long dance generation; however, their settings and corresponding models are not released. Therefore, we do not include comparisons with those methods.
@@
Method 7.5 seconds 10 seconds $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ EDGE [9] 0.26 1.11 59.43 25.60 2.97 3.52 0.25 0.88 68.63 32.24 2.55 3.14 SGMD (ours) 0.30 1.38 55.65 20.74 4.10 5.74 0.30 2.13 56.15 29.98 5.21 6.45 Ground Truth 0.38 1.04 9.34 7.47 0.49 1.70 9.34 7.47
IV-C Settings of Controllable Dance Generation
As there are no relevant benchmarks, we set up experiments for controllable dance generation by integrating spatial constraints. The settings for representative controllable generation are as follows.
Trajectory: Given a predefined trajectory, the model generates dance movements that are coherent with the trajectory.
@@ IV-C Settings of Controllable Dance Generation
Seed Motion: Given motions of the first and second frames, the model generates motions of the following frames.
@@
In-betweening: Given motions of the first and last frames, the model completes motions of the in-between frames.
@@
TABLE III: Results of controllable dance generation given different spatial constraints.
Conditions Method $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ Trajectory EDGE [9] 0.27 1.51 31.1 16.23 5.03 5.23 SGMD (ours) 0.33 1.23 32.22 14.85 5.19 6.02 Seed motion EDGE [9] 0.27 1.19 32.62 21.68 5.81 5.15 SGMD (ours) 0.33 1.18 27.05 17.36 7.34 5.79 In-betweening EDGE [9] 0.28 1.17 37.33 15.16 4.58 4.86 SGMD (ours) 0.34 1.24 30.20 14.75 5.80 6.16 Inpainting EDGE [9] 0.36 2.15 21.57 17.50 7.47 8.36 SGMD (ours) 0.37 1.71 23.80 18.78 7.03 8.34 Upper-body generation EDGE [9] 0.29 1.67 32.90 14.82 5.05 5.45 SGMD (ours) 0.33 1.40 32.86 14.36 5.39 6.11 Lower-body generation EDGE [9] 0.33 0.92 39.81 15.44 5.28 6.86 SGMD (ours) 0.36 0.99 42.26 14.73 5.39 7.37 Ground Truth 0.35 1.33 9.43 7.33
IV-D Results of Dance Generation
We choose EDGE [9] as our baseline. Since the reported results are averaged over 100 trials to reduce randomness, we re-implement the results of EDGE and FACT on the AIST++ dataset. For different approaches, we generate 20 pieces of dances with 5-second length, covering 10 different genres. Quantitative results are summarized in Table I. When comparing our approach with state-of-the-art methods, we observe a significant improvement in the Beat Alignment Score, the primary metric for dance generation, suggesting that the dances generated of our method are more in tune with the musical melody. Specifically, our method performs favorably against EDGE on most metrics. Our method achieves the lowest $FID_{k}$ among diffusion-based models. This metric, defined based on motion velocities and energies, reflects the physical characteristics of dance. The superior $FID_{k}$ score demonstrates that the proposed SGMD generates more realistic dance movements. Our approach also achieves the closest approximation to the ground-truth values of $Div_{k}$ and $Div_{g}$ among diffusion-based models, demonstrating its ability to generate more diverse dance movements instead of converging to a limited set of fixed future motions.
IV-E Results of Long-Term Dance Generation
Since both the training and test sets are composed of five-second slices, in order to validate the ability of our model to generate long-term dance sequences, we generate the corresponding dances for music that are 7.5 seconds and 10 seconds. We select the pieces of dances that are longer than 7.5 seconds or 10 seconds for long-term dance generation. Experimental result of long-term dance generation on AIST++ is shown in Table II. Since there is limited work on long dance generation, we compare our approach only with the baseline.
Our method performs significantly better than EDGE [9] on most metrics. Specifically, our model achieves a higher Beat Alignment Score, maintaining good performance on the correlation between dance movements and music. Our model also performs better than EDGE on metrics of $FID_{k}$, $FID_{g}$, $Div_{k}$ and $Div_{g}$. This reveals that SGMD can also generate long dance movements with better quality and better diversity. It is worth noting that the performance of our model on $Div_{k}$ and $Div_{g}$ do not decrease significantly over time like EDGE. This suggests that long dance movements generated using our method can ensure diversity, avoiding repetitive movements.
IV-F Results of Controllable Dance Generation
Our approach is applicable to controllable dance generation given different spatial constraints. We quantitatively compare our approach with EDGE [9] in Table III.
Given the predefined trajectory, our method outperforms EDGE in terms of the Beat Alignment Score by 0.06, representing a significant relative improvement of 22%. Given conditions of both the seed motion and the upper-body, our method performs favorably against EDGE [9] on all metrics. For dance in-betweening, dance inpainting and lower-body generation, our method performs better than EDGE [9] on most metrics. This suggests that our approach has a strong capacity for flexible generation from both the temporal and spatial perspectives. Furthermore, it is found that the performance of our model on Beat Alignment Score is close to the ground truth or even exceeds it, revealing the potential for self-improvement properties in harmonising music and movement. Examples of controllable dance generation given a predefined trajectory are shown in Figure 4. Overall, our approach is suitable for a wide range of tasks of controllable dance generation.
Figure 4: Illustrations of music-conditioned controllable dance generation given a predefined spatial trajectory.
TABLE IV: Comparison of results with different style prompts.
Style Prompts $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ One-hot Encoding 0.29 1.61 40.13 19.18 Genre Name 0.26 1.96 32.36 17.54 Style Description 0.31 1.51 34.32 19.04
TABLE V: The impact of different audio representations.
@@ IV-F Results of Controllable Dance Generation
Method $BeatAlign$ $PFC$ $FID_{k}$ $FID_{g}$ $Div_{k}$ $Div_{g}$uparrow $\downarrow$ $\downarrow$ $\downarrow$ $\rightarrow$ $\rightarrow$ Jukebox 0.31 1.51 34.32 19.04 6.29 5.81 Encodec 0.28 1.36 43.67 18.97 4.41 5.08 Librosa 0.29 1.48 32.78 16.60 5.80 6.54
TABLE VI: Ablation studies of the proposed SGMD. For simplicity, style modulation and style description prompt are shortened to modulation and description, respectively.
Method $BeatAlign$ $PFC$ $FID_{k}$ $FID_{g}$ $Div_{k}$ $Div_{g}$uparrow $\downarrow$ $\downarrow$ $\downarrow$ $\rightarrow$ $\rightarrow$ Baseline 0.26 1.56 35.35 18.92 5.32 4.90 w/ Modulation 0.26 1.96 32.36 17.54 6.46 5.22 w/ Description 0.29 1.71 31.57 18.47 5.84 5.30 Ours 0.31 1.51 34.32 19.04 6.29 5.81
IV-G Ablation Studies and Discussions
We conduct detailed ablation studies and analyze the impacts of various audio representations and style prompts.
Comparisons of Different Style Prompts: In Table IV, we compare three types of style prompts: one-hot encoding, genre name and style description. We find that using style description prompts performs better than the other two prompts on most metrics. Compared to one-hot encoding and genre name, style description prompts contain richer semantic information related to musical styles, which enhances the quality of dance generation. Since the textual style description is only used for Style Modulation, which is lightweight, the computational cost introduced by this module can be ignored.
Comparisons of Different Audio Representation: In Table V, we evaluate the impact of three independent music feature extractors: Jukebox [39], Encodec [40], and Librosa [41]. Jukebox, a pre-trained generative model designed for music generation, showcasing notable performance in tasks specific to music prediction as evidenced in prior studies [9]. Encodec is a pre-trained model, which has been used as music conditional feature extractor in dance generation tasks as well. Librosa is a library for analyzing audio signals. We observe that the Beat Alignment Score and $Div_{k}$ achieve the optimal performance when Jukebox is utilized as the feature extractor; PFC attains its peak performance with Encodec serving as the feature extractor; and $FID_{k}$, $FID_{g}$ and $Div_{g}$ exhibit the best results when Librosa is employed as the feature extractor. Since the Beat Alignment Score is the primary metric, we use Jukebox to extract audio features.
Effectiveness of Style Modulation and Description Prompts: Ablation studies of the proposed SGMD are shown in Table VI. To verify the effectiveness of the Style Modulation module, we remove the Style Modulation module and simply replace it with concatenating music conditions $c$ and style description prompts $s$, which then enters the Transformer-based block. We find that this variant, referred to as w/ modulation, does not improve the Beat Alignment Score over the baseline.
In addition, we use genres instead of style description prompts, and feed its encoding to the style modulation. This approach, denotes as w/ description, also performs worse on most metrics compared to ours. The possible reason for this is that genres contain less information than style description prompts. To investigate the importance of combining Style Modulation and Style Description Prompts, we remove the Style Modulation module and omit the use of any prompts, which corresponds to the baseline. We can see that the performances are obviously worse in this case. We can conclude that it is quite important for Style Modulation and Style Description Prompts to work together.
The beat alignment score calculates the average time distance between each music beat and its closest dance beat. However, dancing is not solely about syncing local speed minima in joints with beats. Instead, musical beats serve as a flexible reference for timing, rhythm, and the smooth transition between dance movements and steps. While this metric has driven progress on this issue in the past, more dance-specific metrics are needed to realize breakthroughs.